OpenStudy (anonymous):

(Fill in box) A CD makes 58 1/2 revolutions in 3/4 min. Complete the unit rate. [ ] revolutions/min

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Rewrite 58 1/2 as a whole number ratio (fraction) \(\frac{58 \cdot 2 +1}{2} = \ ?\)

OpenStudy (anonymous):

57/2?????

OpenStudy (jhannybean):

(58 x 2) +1 = ?

OpenStudy (anonymous):

117

OpenStudy (jhannybean):

So it's \(\frac{117}{2}\). Now we can write this as a decimal : 58.5 rev. Net we have \(\frac{3}{4}\). Is you plug this into your calculator, what decimal value do you get?

OpenStudy (anonymous):

????????????

OpenStudy (jhannybean):

You know that 1/4 can be written as 0.25, yes?

OpenStudy (anonymous):

yes :)

OpenStudy (jhannybean):

Think of a dollar. It has 4 quarters. If 4 quarters = 1 dollar, then 1 quarter/4 quarters = 25 cents. So what does 3 quarters/4 quarters = ?

OpenStudy (anonymous):

divide

OpenStudy (jhannybean):

No what value do you get.

OpenStudy (anonymous):

3/4 im really bad at this

OpenStudy (jhannybean):

We got 1/4 = 25 cents, so if you have 3 quarters how many do you have?

OpenStudy (anonymous):

75

OpenStudy (jhannybean):

Good, finally.

OpenStudy (jhannybean):

So 3/4 = 0.75 min. Now we just input this into our calculator. 58.5 rev / 0.75 min = ___________ rev/min

OpenStudy (jhannybean):

With this you will find your answer :)

OpenStudy (anonymous):

78??